Sol Summary – January 03rd

  

Sol: 12

Summary Title: Mars Stars

Author's name: Pat Pesa

Mission Status: Sad that our mission is coming to a close.

Sol Activity Summary: Today we had our last EVA ( :( ) and then started the long process of packing and cleaning for the next crew coming tomorrow.

Look Ahead Plan: Tomorrow we look forward to finishing cleaning and welcoming Crew 219!

Anomalies in work: None

Weather: Foggy and cloudy, around 30 degrees

Crew physical status: Healthy

EVA: EVA #12 this morning was our last EVA of the mission. They went north on Cow Dung Road for continued radio signal and weather observations.

Reports to be filed: sol summary, commander report, operations report, greenhab report, journalist report, EVA report, science report

Support Requested: none

Pat Pesa

Geologist, MDRS 218